OpenStudy (anonymous):

Please help, I'll give you a medal<3 You want to pick a four-player tennis team. There are eight players to choose from. How many different teams can you form? A.70 B.1680 C.48 D. 336

jimthompson5910 (jim_thompson5910):

How many people are there to choose from?

OpenStudy (anonymous):

Eight @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Once you made that selection for the first slot, you have 8-1 = 7 choices for slot 2 Then you'll have 8-2 = 6 choices for slot 3 and finally 8-3 = 5 choices for slot 4

jimthompson5910 (jim_thompson5910):

Multiply out those choices: 8*7*6*5 = 1,680 So there are 1,680 ways to pick 4 people if order mattered

jimthompson5910 (jim_thompson5910):

If order doesn't matter, then there are 1,680/(4*3*2*1) = 1,680/24 = 70 ways to do it